100 ml of water requires 16.8 molecule M/50 disodium EDTA for end point
in titration. 100 ml of the same water sample after boiling and filtration
takes 10.7 ml of the disodium EDTA for end point in titration. Calculate
Temporary & permanent hardness of water sample in CaCO3 equivalent.
1
Expert's answer
2020-10-26T14:48:19-0400

Solution.

"H(Temporary) = \frac{C(EDTA) \times V(EDTA) \times 1000}{V(sample)}"

"H(Temporary) = \frac{0.02 \times 16.8 \times 1000}{100} = 3.36 mmol-equiv/L"

"H(Permanent) = \frac{C1(EDTA) \times V1(EDTA) - C2(EDTA) \times V2(EDTA) \times 1000}{V(sample)}"

"H(Permanent) = \frac{(0.02 \times 16.8-0.02 \times 10.7) \times 1000}{100} = 1.22 mmol-euiv/L"

Answer:

H(Temporary) = 3.36 mmol-equiv/L

H(Permanent) = 1.22 mmol-euiv/L